OpenStudy (anonymous):

Which of the following is the solution of log ( base x – 3) 0.001 = –3 ? x = 15 x = 5 x = 7 x = 13 Which of the following is the solution of log ( base x – 3) 0.001 = –3 ? x = 15 x = 5 x = 7 x = 13 @Mathematics

OpenStudy (anonymous):

\[\log_{x-3}(.001)=-3 \] First exponentiation both sides to remove the "log" \[.001=(x-3)^{-3}\] invert both sides to get rid of the negative exponent \[1/.001=1000=(x-3)^{3}\] take the cube root of both sides \[10=x-3\]